Part A What magnetic field B is needed to keep 999-GeV protons revolving in a circle of radius 2.0 km ? Use the relativistic mass. The protons rest mass is 0.938 GeV/c2 1 GeV - 10 eV) [Hint: In relativity, mrelvr-quB in still valid in a magnetic field, mrel-m Express your answer to two significant figures and include the appropriate units B-Value Units Submit Request Answer
Answer #1

Energy of proton,

E = KE + mpc^2

E = 999 + 0.938 = 999.938 GeV

Relative mass, mrel= \gamma mp

999.938 = \gamma * 0.938

\gamma = 1066.03

From equilibrium of force,

qvB = mrel * v^2 / r

B = mrel * v / qr = \gamma *mp v / qr

B = 1066.03 * 1.67*10-27 * 3*108 / 1.6*10-19 * 2*103

B = 1.669 T

Magnetic field B = 1.669 T